[Brahms-l] Shift report 20040313 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sat Mar 13 2004 - 16:00:56 EST
------------------------------------------------------------
Date: Sat Mar 13 16:00:55 2004, Shift: 08:00-16:00
Supervisor: hiro, with: 
Log book#15 pages 86 - 86

Shift Summary:
------------------------------------------------------------
1.  Beam for the most part.
    Quench link interlock.
2.  finish FS 10deg A608
    Finsh MRS zero
    Finish MRS 1050A
    Finish 350A.
3.  DAQ run server problem continues. I tried several fix without any good results. 

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Mar 13 16:01:13 2004

This archive was generated by hypermail 2.1.8 : Sat Mar 13 2004 - 16:01:34 EST